# Cosmic dynamics and observational constraints in $f(Q)$ gravity with affine equation of state

### v1

@@ 4 Observational constraints and results
Within this part, we focus on describing the data sets that are utilized to analyze the models. We also explain constraints on Model I(15) and Model II(16)’s model parameters. In other words, the consistency of the Hubble parameter (15) and (16) with the data sets given by the Cosmic Chronometer (CC) and combined data that consists of Pantheon and CC are used in the MCMC analysis. We constrain the parameters ($H_{0}$, $\alpha$, $\rho_{0}$ and $\Omega_{m0}$) using the $\chi^{2}$ minimization method and the MCMC method is implemented with the emcee Python library [75].
@@ 4.1 The Cosmic chronometer data
The model parameters ($H_{0}$, $\alpha$, $\rho_{0}$ and $\Omega_{m0}$) are constrained using Cosmic Chronometer data consisting of $31$ data points in the red-shift range of $0.07\leq z\leq 1.965$ are collected using the differential ages of galaxies approach [76, 77]. The cosmic chronometer observations’s fundamental principle (introduced by Jimenez and Loeb [78]) introduces a relationship between red-shift $(z)$, cosmic time $t$, and the Hubble parameter $H(z)$ as $H(z)=\frac{-1}{(1+z)}\frac{dz}{dt}$. By minimizing the $\chi^{2}$ function, we determine the model parameters’s central values with $1\sigma$ errors in the emcee[75] package. For the Cosmic chronometer data, the $\chi^{2}$ function may be represented as [71, 79, 80, 81]:
@@
Here, $H_{th}$ is used to indicate the theoretical values of the Hubble parameter, the observed values are indicated by $H_{obs}(z_{i})$ and the notation $\sigma_{i}$ indicates the standard deviation for each $H_{obs}(z_{i})$ observed value.
@@
Figure 1: In comparison to the $\Lambda CDM$ model, the best fit Hubble parameter curves versus $\mathit{z}$.
It may observed that the cosmic dynamics of the Model I and II are deviating from the $\Lambda$CDM model in the past where $z>1$. The transition redshift however remains close to the $\Lambda$CDM model. It is worthwhile to observe from these results that the Model I may describe the transiting universe evolution from the decelerated expansion into the accelerated expansion and the even with the consideration of pressure-less matter in Model II, the cosmic dynamics remains identical as that of the Model I.
@@ 4.2 The Pantheon data
${}m_{th}(z)=M+5\log_{10}\left[\frac{d_{L}(z)}{1Mpc}\right]+25,$ (18)
@@
${}m_{th}(z)=M+5\log_{10}\left[D_{L}(z)\right]+5\log_{10}\left[\frac{c/H_{0}}{1Mpc}\right]+25.$ (19)
In the $\Lambda$CDM model framework, a degeneracy will exist between $H_{0}$ and $M$ [83, 84]. We take $\mathcal{M}$ as a combination of these parameters as [84]:
${}\mathcal{M}\equiv M+5\log_{10}\left[\frac{c/H_{0}}{Mpc}\right]+25$ (20)
where $H_{0}=h\times 100$ Km/(s.Mpc) and thus one can write $\mathcal{M}=42.39+M-5\log_{10}(h)$. For the Pantheon data, this parameter is used along with the relevant $\chi^{2}$ in the MCMC analysis defined as [84, 80, 79, 71, 81]
@@
where, $\nabla m_{i}=m_{obs}(z_{i})-m_{th}(z_{i})$, the theoretical apparent magnitude $m_{th}(z_{i})$ is calculated using equation (19) and the symbol $m_{obs}(z_{i})$ is used to denote the apparent magnitude of SNeIa for red-shift $z_{i}$ and $C_{ij}^{-1}$ is covariance matrix’s inverse. The luminosity distance is influenced by the Hubble parameter. Following this approach, we use equations (15) and (16) and the emcee package [75] to determine the maximum likelihood estimate using the joint (CC+Pantheon) dataset. For computing the maximum likelihood estimate using the joint data, we define the joint $\chi^{2}$ as $\chi^{2}_{CC}+\chi^{2}_{P}$.
@@
Table 1: Model I: For CC and joint data, the constrained values along with the values of $q_{0},z_{t}$, $t_{0}$ $j_{0}$.
@@
CC $68.23^{+0.69}_{-0.69}$ $0.024^{+0.015}_{-0.024}$ $4.67^{+1.3}_{-0.72}$ $0.157^{+0.036}_{-0.030}$0.53 0.64 $13.58^{+0.33}_{-0.22}$ $1.016$ $-0.690$
joint $68.6^{+1.8}_{-1.8}$ $0.025^{+0.014}_{-0.014}$ $4.76^{+1.2}_{-0.51}$ $0.150^{+0.023}_{-0.039}$ $23.81^{+0.0099}_{-0.0099}$0.54 0.64 $13.59^{+0.69}_{-0.32}$ $1.017$ $-0.698$
Table 2: Model II: For CC and joint data, the constrained values along with the values of $q_{0},z_{t}$, $t_{0}$ and $j_{0}$.
Figure 2: Model I: Marginalized $1\sigma$ and $2\sigma$ contour map with median values of $H_{0}$, $\alpha$ and $\rho_{0}$ using Joint data set.
@@ 5.1 Cosmographic parameter
The study of cosmography was initially introduced by Weinberg [85], who used a Taylor series to introduced the scale factor that rose around present time $t_{0}$. A detailed overview of cosmography may be reviewed for different aspects of cosmographic parameters [23]. The Hubble parameter ($H$), has been considered as a variable observable quantity. The behavior of Hubble parameter (1) highlights that the expansion rate of the universe in considered models. The evolution of deceleration parameter $q$ is characterized through the second-order derivative of scale factor ($a$) [86]. To understand the universe’s cosmic evolution, the snap ($s$) and jerk ($j$) parameters serve as important tools. One of the cosmological parameter is the deceleration parameter that helps to characterize the universe’s expansion behavior. It is basically used to track down the rate of slowing down of the universe expansion. Mathematically, using the Hubble parameter and its derivative, it may be expressed as
@@
This parameter serves as a tool to assess whether the universe is expanding at an increasing or decreasing rate. Depending on the value of $q$, different phases of cosmic expansion may be visualized. If $q$ is greater than zero $(q>0)$, the universe is in a decelerated phase. For $q<0$, the universe will expanding at an accelerating rate. For $-1<q<0$ the expansion follows a power-law acceleration. A super-exponential expansion era occurs in the universe for $q<-1$, while a de Sitter expansion when $q=-1$ [87, 88, 89, 90]. For matter and radiation-dominated universe expansion, the deceleration parameter will be $q=\frac{1}{2}$ and $q=1$ respectively. By using (15), (16) and (22), we may obtain for Model I and II as
$q(z)=-1+\left({\frac{(1.5)(1+\alpha)(1+z)^{3+3\alpha}}{\rho_{0}+(1+z)^{3+3\alpha}}}\right)\ \ \text{and}$ (23)
$q(z)=-1+\frac{0.5(1+z)^{3}\left(3\Omega_{m0}+\left(\frac{3(1-\Omega_{m0})(1+z)^{3\alpha}(1+\alpha)}{(1+\rho_{0})}\right)\right)}{\left(\Omega_{m0}(1+z)^{3}+\frac{(1-\Omega_{m0})(\rho_{0}+(1+z)^{3+3\alpha})}{(1+\rho_{0})}\right)}$ (24)
@@
Figure 3: Model I: The deceleration parameter with $z$.
Figure 4: Model II: The deceleration parameter with $z$.
@@
The deceleration parameter of these models provides insight into the universe’s transition from deceleration into acceleration of expansion. In model I, the present day value of deceleration parameter is $q=-0.56$ and $q=-0.58$ for CC and joint data respectively, and the value of transit red-shift are $z_{t}=0.638$ (for CC data) and $z_{t}=0.65$ (for joint data). The red-shift value at which $q=0$ is termed as the transit red-shift, since the behavior of $q$ is changing from positive to negative in the cosmic evolution. The accelerated universe expansion may be visualized in the model I for red-shifts $z<0.638$. For Model II, the transition from decelerated to accelerated expansion occur at $z_{t}=0.64$ based on CC estimates and at $z_{t}=0.647$ using joint estimates. For CC estimates, the deceleration parameter’s current value is $q_{0}=-0.536$, while for joint estimates, it is $q_{0}=-0.548$. These negative value of deceleration parameter of model I and II are confirming that the universe is expanding with acceleration in the present era $(z=0)$. And, thus it has been observed that the non-metricity scalar may yield the transiting universe evolution in model I governed by $f(Q)=C\sqrt{Q}$. In the Model II, the dark energy introduced by the non-metricity may yield the accelerated expansion at late-times while the cold dark matter yields the decelerated universe expansion at early times in model.
The behavior of $q$ in Fig. (4) and (4) suggests that these models will possess the matter-dominated era of universe expansion which may be confirmed with $(q\rightarrow\frac{1}{2})$ during early times. An accelerated phase $(q<0)$ in the present era and ultimately converges to the de Sitter scenario $(q=-1)$ as $z\rightarrow-1$. These models may explain the accelerated expansion universe at the present times, which is consistent with the observational data. It has been observed that the deceleration parameter of Model I and Model II are consistent with the current phase of accelerated expansion.
The jerk parameter $(j)$ indicates the rate at which the universe’s acceleration or deceleration changes over time. The snap parameter $(s)$ calculates the rate that the jerk parameter changes. These parameters are defined as follows:
@@
In the present analysis, for $j$ and $s$, equation (25) can be rewritten in terms of red-shift [91].
@@
For Model I, based on the constrained values from the CC data set, the snap parameter is $s_{0}=-2.12$ and the jerk parameter $j_{0}=1.14$, however, from the joint estimates, the jerk parameter value is $j_{0}=1.25$ and snap parameter value is $s_{0}=-2.03$. Similarly, for Model II, the value of jerk parameter is $j_{0}=1.0164$ (for CC data set) and $j_{0}=1.017$ (for joint data set). In the Tables (1) and (2), we summarize the present day values of the cosmological parameters. For Models I and II, the value of jerk parameter exhibits a decreasing trend from early to late times, eventually approaching to $1$. This behavior suggests that these models differs from $\Lambda$CDM in the early universe but becomes similar to $\Lambda$CDM at later times.
5.2 The physical behavior of models based on the energy density, pressure and equation of state parameter
@@
We highlight the physical behavior exhibited by the quantities such as the energy density and pressure. For the model parameters’s constrained values, the energy density has remained positive throughout the expansion history while the pressure may have undergone a recent transition to negative values. We calculate the energy density and pressure by utilizing equations (9), (15) and (16) as follows:
$\rho_{d}(z)=3H_{0}^{2}\left[\frac{(1+z)^{3(1+\alpha)}+\rho_{0}}{\rho_{0}+1}\right]\quad\text{for Model I}{}$ (27)
$p_{d}(z)=\frac{3H_{0}^{2}(1+\alpha)(1+z)^{3+3\alpha}}{(\rho_{0}+1)}-3H_{0}^{2}\left[\frac{(1+z)^{3(1+\alpha)}+\rho_{0}}{\rho_{0}+1}\right]\quad\text{ for Model I}$ (28)
$\rho_{e}(z)=3H_{0}^{2}\left[\Omega_{m0}(1+z)^{3}+\left[\frac{1-\Omega_{m0}}{\rho_{0}+1}\right]\left[(1+z)^{3(1+\alpha)}+\rho_{0}\right]\right]\quad\quad\text{for Model II}$ (29)
$p_{e}(z)=3H_{0}^{2}(1+z)^{3}\left[\Omega_{m0}+\frac{(1-\Omega_{m0})(1+\alpha)(1+z)^{3\alpha}}{(1+\rho_{0})}\right]-3H_{0}^{2}\left[\Omega_{m0}(1+z)^{3}+\left[\frac{1-\Omega_{m0}}{\rho_{0}+1}\right]\left[(1+z)^{3(1+\alpha)}+\rho_{0}\right]\right]\quad\text{for Model II}$ (30)
@@
The evolution of energy density is depicted in figures (6) and figure (6) for Model I and II respectively. These figures are revealing that the energy densities in the models will be increasing with red-shift $(z)$ (decreases with cosmic time $(t)$), remaining positive throughout expansion for both the models. In other words, the energy densities will be decreasing from past to future. The energy densities will have positive values rules out the existence of finite-time future singularities in the models. During the early universe, at large red-shift ($z$), the pressure commences with positive values. In the later epoch along with the present epoch, the pressure becomes negative. It is observed from the graphical analysis that the non-metricity scalar will be responsible for the negative pressure necessary for the accelerated universe expansion in this late epoch.
Figure 5: Model I: $\rho_{d}$ versus $\mathit{z}$
Figure 6: Model II: $\rho_{e}$ versus $\mathit{z}$
@@
The EoS parameter may offer insights on the dominating fluid during the cosmic expansion of universe and thus helps to determine the dynamics of Universe. In the cosmological models, it may be commonly used to identify the nature of dark energy. The EoS parameter can be defined using the relationship between pressure $(p)$ and energy density $(\rho)$. It is mathematically expressed as $(\omega=\frac{p}{\rho})$. The cosmic characteristics, including the cold dark matter or dust dominated phase will have $\omega=0$ and, the radiation-dominated period will have $\omega=\frac{1}{3}$. The transition into an accelerated cosmic era happens for the equation of state parameter satisfying $\omega<-\frac{1}{3}$. The cosmic acceleration period may have three regimes: the quintessence-dominated regime for $-1<\omega<-\frac{1}{3}$, the cosmological constant-like state for $\omega=-1$ and phantom-dominated regime $\omega<-1$. In the Figs. (8) and (8), we provide the graphical representation of the EoS parameter evolution in the considered models.
Figure 7: Model I: $\omega_{d}$ versus $z$
Figure 8: Model II: $\omega_{e}$ versus $z$
In the Model I, the EoS parameter values are estimated as $\omega_{d}=-0.706$ and $\omega_{d}=-0.720$ for CC and joint dataset respectively at $z=0$. Similarly, in the Model II, the effective EoS parameter values are obtained $\omega_{e}=-0.690$ (for CC dataset) and $\omega_{e}=-0.698$ (for joint dataset) at $z=0$. The Fig. (8) and (8) are revealing that the EoS parameter may correspond to the quintessence kind of dark energy in these models. At late-times, as $z\rightarrow-1$, the cosmological constant-like behavior may be visualized in these models. In the present era, the dark energy is a dynamical one which is varying with time in these models.
@@
These models possess the past of matter-dominated expansion which are confirmed by Fig. (8) and (8) having $\omega=0$ during early times. The cosmic evolution is under the domination of quintessence kind of dark energy having $-1<\omega<-\frac{1}{3}$ in the present epoch, and ultimately converges towards the $\Lambda$CDM scenario $\omega=-1$ as $z\rightarrow-1$. For these evolution scenarios, we observe the role of non-metricity scalar having geometrical origin in the $f(Q)$ model. In the model I, the non-metricity is in charge for the matter dominated expansion and the dark energy dominated expansion. In this sense, the unified scenario has been confirmed in the affine EoS induced model of $f(Q)$ gravity. On the other hand, when the cold dark matter has been considered with the dark fluid satisfying affine EoS, the $f(Q)$ gravity model yields the lower $\Omega_{m0}$ value as compared to the $\Lambda$CDM model. The possible reasons for this change may be the matter-like characteristics possessed in the past by dark fluid having affine EoS. This is an important difference between the cosmic dynamics possessed by model without and with cold dark matter. In the model without dark matter, the fluid with affine EoS unifies the decelerated and accelerated expansion. While in the model with cold dark matter and fluid having affine EoS, the matter dominated era (having $q=\frac{1}{2}$) is controlled by both of these components leading to the $\omega_{e}=0$ scenario in effective sense.

### v2

@@ 4 Observational constraints and results
In this section, we focus on describing the data sets and methodology that are utilized to analyze the models. We also explain constraints on Model I (15) and Model II (16)’s model parameters. In other words, the consistency of the Hubble parameter (15) and (16) with the data sets given by the Cosmic Chronometer (CC) and combined data (consisting of Pantheon and CC) are used in the MCMC analysis. We constrain the parameters using $\chi^{2}$ minimization method and the MCMC method is implemented with the emcee Python library [79].
@@ 4.1 The Cosmic chronometer data
The model parameters are constrained using the Cosmic Chronometer (CC) data consisting of $31$ data points in the red-shift range of $0.07\leq z\leq 1.965$. These data are collected using the differential ages of galaxies approach [80, 81]. The cosmic chronometer observations are based on the principle introduced by Jimenez and Loeb [82]. It describes a relationship between red-shift $(z)$, cosmic time $t$, and the Hubble parameter $H(z)$ as $H(z)=\frac{-1}{(1+z)}\frac{dz}{dt}$. By minimizing the $\chi^{2}$ function, we determine the model parameters’ central values with $1\sigma$ errors in the emcee [79] package. The $\chi^{2}$ function has been taken as [75, 83, 84, 85]
@@
Figure 1: The best fit Hubble parameter curves with $z$ as compared to the $\Lambda CDM$ model.
It may be observed that the cosmic dynamics of the Model I and II are deviating from the $\Lambda$CDM model in the past where $z>1$. The transition redshift however remains close to the $\Lambda$CDM model. It is worthwhile to observe from these results that the Model I may describe the transiting universe evolution from the decelerated expansion into the accelerated expansion and even with the consideration of pressure-less matter in Model II, the cosmic dynamics remains similar as that of the Model I.
@@ 4.2 The Pantheon data
${}m_{th}(z)=M+5\log_{10}\left[\frac{d_{L}(z)}{\text{Mpc}}\right]+25,$ (18)
@@
${}m_{th}(z)=M+5\log_{10}\left[D_{L}(z)\right]+5\log_{10}\left[\frac{c/H_{0}}{\text{Mpc}}\right]+25.$ (19)
In the $\Lambda$CDM model framework, a degeneracy will exist between $H_{0}$ and $M$ [87, 88]. We take $\mathcal{M}$ as a combination of these parameters [88]:
${}\mathcal{M}\equiv M+5\log_{10}\left[\frac{c/H_{0}}{\text{Mpc}}\right]+25$ (20)
where $H_{0}=h\times 100\ Km/s/Mpc$ and thus one can write $\mathcal{M}=42.39+M-5\log_{10}(h)$. For the Pantheon data, this parameter is used along with the relevant $\chi^{2}$ in the MCMC analysis defined as [88, 84, 83, 75, 85]
@@
where, $\nabla m_{i}=m_{obs}(z_{i})-m_{th}(z_{i})$, the theoretical apparent magnitude $m_{th}(z_{i})$ is calculated using equation (19) and the symbol $m_{obs}(z_{i})$ is used to denote the apparent magnitude of SNeIa observations at red-shift $z_{i}$ and $C_{ij}^{-1}$ is covariance matrix’s inverse. The luminosity distance is influenced by the Hubble parameter. Following this approach, we use equations (15) and (16) and the emcee package [79] to determine the maximum likelihood estimate using the joint (CC+Pantheon) dataset. For computing the maximum likelihood estimate using the joint data, we define the joint $\chi^{2}$ as $\chi^{2}_{CC}+\chi^{2}_{P}$.
@@
Table 1: Model I: The constrained values with $q_{0},z_{t}$, $t_{0}$ $j_{0}$, where $H_{0}$ is in $Km/s/Mpc$.
@@
CC $71.0^{+5.5}_{-8.2}$ $0.305^{+0.067}_{-0.056}$ $-1.33^{+0.73}_{-0.49}$ $-0.881$ $0.62$ $2.352$ $13.76$ $-0.921$
joint $68.7^{+1.9}_{-1.9}$ $0.322^{+0.051}_{-0.042}$ $-1.08^{+0.17}_{-0.13}$ $23.806\pm 0.014$ $-0.597$ $0.63$ $1.259$ $13.59$ $-0.7314$
Table 2: Model II: The constrained values with $q_{0},z_{t}$, $t_{0}$ and $j_{0}$, where $H_{0}$ is in $Km/s/Mpc$.
Figure 2: Model I: Marginalized $1$D and $2$D posterior contour map with median values of parameters using Joint data set.
Figure 3: Model II: Marginalized $1$D and $2$D posterior contour map with median values of parameters using Joint data set.
@@ 5.1 Cosmographic parameter
The study of cosmography was initially introduced by Weinberg [89], who used a Taylor series to introduce the scale factor that rose around present time $t_{0}$. A detailed overview of cosmography may be reviewed for different aspects of cosmographic parameters [23]. The Hubble parameter ($H$), has been considered as a variable observable quantity. The behavior of the Hubble parameter (Fig. 1) highlights the expansion rate of the universe in considered models. The evolution of deceleration parameter $q$ is characterized through the second-order derivative of scale factor ($a$) [90]. To understand the universe’s cosmic evolution, the snap ($s$) and jerk ($j$) parameters serve as important tools. One of the cosmological parameters is the deceleration parameter that helps to characterize the universe’s expansion behavior. It is basically used to track down the rate of slowing down of the universe’s expansion. Mathematically, using the Hubble parameter and its derivative, it may be expressed as
@@
This parameter serves as a tool to assess whether the universe is expanding at an increasing or decreasing rate. Depending on the value of $q$, different phases of cosmic expansion may be visualized. If $q$ is greater than zero $(q>0)$, the universe is in a decelerated phase. For $q<0$, the universe will expand at an accelerating rate. For $-1<q<0$ the expansion follows a power-law acceleration. A super-exponential expansion era occurs in the universe for $q<-1$, while a de Sitter expansion when $q=-1$ [91, 92, 93, 94]. For matter and radiation-dominated universe expansion, the deceleration parameter will be $q=\frac{1}{2}$ and $q=1$ respectively. By using (15), (16) and (22), we may obtain for Model I and II as
$q(z)=-1+\left({\frac{3(1+\alpha)(1+z)^{3+3\alpha}}{2(\rho_{0}+(1+z)^{3+3\alpha})}}\right),\ \ \text{and}$ (23)
$q(z)=-1+\frac{3(z+1)^{3}((\alpha+1)\Omega_{a0}(z+1)^{3\alpha}+\Omega_{m0})}{2((z+1)^{3}(\Omega_{a0}(z+1)^{3\alpha}+\Omega_{m0})+\Omega_{\Lambda 0})},$ (24)
@@
Figure 4: $q(z)$ versus $z$.
@@
The deceleration parameter of these models provides insight into the universe’s transition from deceleration into acceleration of expansion. In model I, the present day value of deceleration parameter is $q_{0}=-0.56$ and $q_{0}=-0.58$ for CC and joint data respectively, and the value of transit red-shift are $z_{t}=0.638$ (for CC data) and $z_{t}=0.65$ (for joint data). The red-shift value at which $q=0$ is termed as the transit red-shift, since the behavior of $q$ is changing from positive to negative in the cosmic evolution. The accelerated universe expansion may be visualized in the model I for red-shifts $z<0.638$. For Model II, the transition from decelerated to accelerated expansion occurs at $z_{t}=0.62$ based on CC estimates and at $z_{t}=0.63$ using joint estimates. For CC estimates, the deceleration parameter’s current value is $q_{0}=-0.881$, while for joint estimates, it is $q_{0}=-0.597$. These negative values of deceleration parameters of model I and II are confirming that the universe is expanding with acceleration in the present era $(z=0)$. And, thus it has been observed that the non-metricity scalar may yield the transiting universe evolution in model I governed by $f(Q)=C\sqrt{Q}$. In the Model II, the dark energy introduced by the non-metricity may yield the accelerated expansion at late-times while the cold dark matter yields the decelerated universe expansion at early times in model.
The behavior of $q$ in Fig. (4) suggests that the model I will possess the matter-dominated era of universe expansion which may be confirmed with $(q\rightarrow\frac{1}{2})$ during early times. An accelerated phase $(q<0)$ in the present era and ultimately converges to the de Sitter scenario $(q=-1)$ as $z\rightarrow-1$. This model may explain the accelerated expansion universe at the present times, which is consistent with the observational data. It has been observed that the deceleration parameter of Model II is consistent with the current phase of accelerated expansion. In Fig. (4), the accelerated universe expansion of present times may reach to super-exponential expansion having $q<-1$ in the future. It is an important implication of the Model II, which is classically unstable due to $\alpha<0$.
The jerk parameter $(j)$ indicates the rate at which the universe’s acceleration or deceleration changes over time. The snap parameter $(s)$ calculates the rate at which the jerk parameter changes. These parameters are defined as follows:
@@
In the present analysis, for $j$ and $s$, equation (25) can be rewritten in terms of red-shift as [95].
@@
For Model I, based on the constrained values from the CC data set, the snap parameter is $s_{0}=-2.12$ and the jerk parameter $j_{0}=1.14$. However, from the joint estimates, the jerk parameter value is $j_{0}=1.25$ and snap parameter value is $s_{0}=-2.03$. Similarly, for Model II, the value of the jerk parameter is $j_{0}=2.35$ (for CC data set) and $j_{0}=1.26$ (for joint data set). In the Tables (1) and (2), we summarize the present day values of the cosmological parameters. For Model I, the value of jerk parameter exhibits a decreasing trend from early to late times, eventually approaching to $1$. This behavior suggests that these models differ from $\Lambda$CDM in the early universe but the Model I would become similar to $\Lambda$CDM at later times.
5.2 The physical behavior of models
@@
We highlight the physical behavior exhibited by the quantities such as the energy density and pressure. For the model parameters’ constrained values, the energy density has remained positive throughout the expansion history while the pressure may have undergone a recent transition to negative values. We calculate the energy density and pressure by utilizing equations (9), (15) and (16). In model I, the dark fluid having energy density $\rho_{d}$ will act as the total energy density of universe in model. And, the Model II possess the matter and the dark energy (due to dark fluid) will have the total energy density $\rho_{e}$.
@@
The evolution of energy densities are depicted in Figs. (6) for Model I and II. The energy densities in these models will be increasing with red-shift $(z)$ (decreases with cosmic time $(t)$), remaining positive throughout expansion for both the models. In other words, the energy density will be decreasing from past to future in Model I. And, thus the energy density will have positive values and it rule out the existence of finite-time future singularities. In Model II, the energy density possess phantom nature. During the early universe, at large red-shift, the pressure commences with the positive values. In the later epoch along with the present epoch, the pressure becomes negative. It is observed from the graphical analysis that the non-metricity scalar will be responsible for the negative pressure necessary for the accelerated universe expansion in this late epoch.
Figure 5: Total energy density versus $\mathit{z}$
Figure 6: $\omega_{e}$ versus $\mathit{z}$
@@
The EoS parameter may offer insights on the dominating fluid during the cosmic expansion of the universe and thus helps to determine the dynamics of the Universe. In the cosmological models, it may be commonly used to identify the nature of dark energy. The EoS parameter can be defined using the relationship between pressure $(p)$ and energy density $(\rho)$. It is mathematically expressed as $(\omega=\frac{p}{\rho})$. The cosmic characteristics, including the cold dark matter or dust dominated phase will have $\omega=0$ and the radiation-dominated period will have $\omega=\frac{1}{3}$. The transition into an accelerated cosmic era happens for the equation of state parameter satisfying $\omega<-\frac{1}{3}$. The cosmic acceleration period may have three regimes: the quintessence-dominated regime for $-1<\omega<-\frac{1}{3}$, the cosmological constant-like state for $\omega=-1$ and phantom-dominated regime $\omega<-1$. In the Fig. (6), we provide the graphical representation of the EoS parameter evolution in the considered models. In the Model I, the EoS parameter (defined as $\omega_{e}=p_{d}/\rho_{d}$) values are estimated as $\omega_{e}=-0.706$ and $\omega_{e}=-0.720$ for CC and joint dataset respectively at $z=0$. The Model II possess matter and dark energy, as a consequence, we define $\omega_{e}=p_{e}/\rho_{e}$, where the total energy density and pressure are denoted by $\rho_{e}$ and $p_{e}$ respectively. In the Model II, the effective EoS parameter values are obtained $\omega_{e}=-0.921$ (for CC dataset) and $\omega_{e}=-0.73$ (for joint dataset) at $z=0$. The Fig. (6) is revealing that the effective EoS parameter may correspond to the quintessence kind of dark energy in these models at present times. In future, as $z\rightarrow-1$, the cosmological constant-like behavior may be visualized in the Model I. On the other hand, in the Model II, the phantom kind of dark energy may evolve during the cosmological evolution. In the present era, the dark energy is a dynamical one which is varying with time in these models.
@@
These models possess the past of matter-dominated expansion which are confirmed by Fig. (6) having $\omega=0$ during early times. The cosmic evolution is under the domination of quintessence kind of dark energy having $-1<\omega<-\frac{1}{3}$ in the present epoch. However, the future evolution in both of these models may be different. For these evolution scenarios, we observe the role of non-metricity scalar having geometrical origin in the $f(Q)$ model. In the model I, the non-metricity is in charge of the matter dominated expansion and the dark energy dominated expansion. In this sense, the unified scenario has been confirmed in the affine EoS induced model of $f(Q)$ gravity. On the other hand, when the cold dark matter has been considered with the dark fluid satisfying affine EoS, the $f(Q)$ gravity model yields the identical matter evolution as of the $\Lambda$CDM model but the dark energy behavior will be clearly different. It is an important difference between the cosmic dynamics possessed by models without and with cold dark matter. In the model without dark matter, the fluid with affine EoS unifies the decelerated and accelerated expansion. While in the model with cold dark matter and fluid having affine EoS, the matter dominated era may be the same but the phantom evolution of dark energy may not be ruled out in $f(Q)$ gravity.